3. Find the Maclaurin series for the functions below. (a) (b) (c) sin 31 (d) - 2r3 - 5x + 4 (f) sin? (3) log 1+1 (h)
Answer #1

Ann. - Mcluain series of a suunction in Writton 1cus = { {co) na g). 1cm = é + 160) = 1 a(n) = en it(o) = -1 12 (1) = enAnn. - Mcluain series of a suunction in Writton 1cus = { {co) na g). 1cm = é + 160) = 1 a(n) = en it(o) = -1 12 (1) = enno Even. inco)= n! enco) = -n! in in n-odd. 1(M) = { anco no 1-nt M²_u3+44 45 . (C) Sin(n) = 1cm) ln) = 3.Com n 1(0) = 0. aenco)=0 & nENULO3. so. elth If § e n. nzo